# Emphasizing Discriminative Features for Dataset Distillation in Complex Scenarios

### v1

@@ 4.3 Ablation Study
We conduct an ablation study to evaluate the impact of EDF’s key components, including the supervision dropout ratio, strategies for discriminative area enhancement, and the frequency of activation map updates. Unless otherwise specified, the following results are all based on ConvNetD5.
@@
EDF introduces two key modules: Discriminative Area Enhancement (DAE) and Common Pattern Dropout (CPD). We conduct an ablation study to assess the contribution of each module independently. The results, presented in Table 4, demonstrate that both DAE and CPD significantly improve the baseline performance. DAE’s biased updates toward high-activation areas using activation-based gradient weights effectively enhance the discriminative features in synthetic images. CPD, on the other hand, mitigates the negative influence of common patterns by filtering out low-loss supervision, ensuring that the synthetic images retain their discriminative properties.
DAE CPD Accuracy(%)
39.2
40.3
41.1
41.8
(a) ImageWoof, IPC10
DAE CPD Accuracy(%)
48.9
49.5
51.2
52.6
(b) ImageMeow, IPC10
DAE CPD Accuracy(%)
65.1
66.2
67.5
68.2
(c) ImageYellow, IPC10
@@
The Discriminative Area Enhancement (DAE) component involves two key factors: the enhancement factor $\beta$ and the threshold for activation maps. Ablation studies (Table 7(a)) show that the best performance is achieved when $\beta$ is between 1 and 2. When $\beta<1$, some discriminative areas are diminished rather than enhanced, as their gradient weights become $<1$. Conversely, excessively large $\beta$ values ($\geq 10$) lead to overemphasis on certain areas, distorting the overall learning process (see Appendix C.3 for examples of this distortion). Therefore, $\beta$ should be reasonably controlled to balance the emphasis on discriminative regions.
Regarding the threshold for activation maps, using the mean activation value as a dynamic threshold results in better performance compared to using a fixed threshold (Table 7(b)). This is because the mean adapts to the evolving activation maps during training, whereas a fixed threshold risks either emphasizing low-activation areas if set too low or omitting key discriminative features if set too high.
@@
Table 7: (a) Ablation of the enhancement factor on ImageMeow and ImageWoof, both IPC10. (b) Ablation of the activation threshold on ImageMeow IPC10.
@@ 5 Analysis and Discussion
Figure 5: (a) Comparison of discriminative areas in images produced by initialization, DATM, and EDF. Figures at the bottom are increments made by EDF over the initial image. (b) Comparison between high-loss and low-loss supervision distilled images.
@@
Our Discriminative Area Enhancement (DAE) module aims to amplify updates in high-activation areas of synthetic images, as identified by Grad-CAM. To show how DAE enhances discriminative areas, we visualize the same group of images under initialization, DATM distillation, and EDF distillation in Figure 5(b) We also report discriminative area statistics, computed by the percentage of pixels whose activation values are higher than the mean, on each image at the bottom left. As can be discovered, DATM is capable of increasing discriminative regions, while EDF can achieve a more significant enhancement. Visually, the enhancement manifests through an increased number of core objects and enlarged areas of class-specific features.
Moreover, EDF’s enhancement is more pronounced especially when the image has smaller discriminative areas initially, e.g. discriminative features of the first column image increase by 24.5%. These phenomena demonstrate that EDF is effective in capturing and emphasizing discriminative features.
@@
To assess the effectiveness of supervision dropout strategies, we compare several dropout approaches. These strategies are classified into two categories: (i) dynamic dropout, which includes random selection from all layers, and (ii) static dropout, which includes uniform selection across layers and fixed selection from the first, middle, or last layers. As shown in Table 8, all strategies except EDF’s loss-based dropout lead to performance degradation, with uniform selection and last-layer dropout causing the most significant performance loss.
The reasons for this are twofold. First, low-loss trajectory parameters—primarily located in the shallow layers of the model—are the main source of common patterns. Discarding supervision from deep layers, where loss values are higher (random selection, uniform selection, or last-layer dropout), reduces the presence of discriminative features. Second, static dropout fails to account for the dynamic nature of low-loss supervision, as trajectory-matching losses vary across layers as the distillation process evolves. By addressing these issues, EDF’s loss-based dropout in CPD mitigates the effects of common patterns and yields superior performance.
@@ 6 Related Work
Dataset Distillation (DD) aims to create compact datasets that maintain performance levels comparable to full-scale datasets. It can be applied in practical fields such as continual learning (Masarczyk & Tautkute, 2020; Rosasco et al., 2021), privacy preservation (Dong et al., 2022; Yu et al., 2023), and neural architecture search (Jin et al., 2018; Pasunuru & Bansal, 2019). Approaches in DD can be categorized into two primary approaches: matching-based and knowledge-distillation-based.
1) Matching-based approaches are foundational in DD research, focusing on aligning synthetic data with real datasets by capturing essential patterns. Landmark works like gradient matching (DC) (Zhao et al., 2021), distribution matching (DM) (Zhao & Bilen, 2021a), and trajectory matching (MTT) (Cazenavette et al., 2022) extract critical metrics from real datasets, then replicate these metrics in synthetic data. Subsequent research has refined these methods, improving the fidelity of distilled datasets (Zhao & Bilen, 2021b; Wang et al., 2022; Zhao et al., 2023; Lee et al., 2022b; Liu et al., 2023a; b; Cazenavette et al., 2023; Sajedi et al., 2023; Khaki et al., 2024). Data selection techniques have been integrated to synthesize more representative samples (Xu et al., 2023; Sundar et al., 2023; Lee & Chung, 2024). Recent advancements optimize distillation for different image-per-class (IPC) settings, balancing dataset size and information retention (Du et al., 2023; Chen et al., 2023; Guo et al., 2024; Li et al., 2024; Lee & Chung, 2024). Moreover, soft labels have been widely applied to improve the performance (Sucholutsky & Schonlau, 2021; Cui et al., 2022a; Qin et al., 2024; Yu et al., 2024). Despite these improvements, most matching-based approaches treat all pixels uniformly, failing to emphasize discriminative regions and often overlooking distinctions between supervision signals, limiting their effectiveness on complex datasets like ImageNet-1K.
2) Knowledge-distillation-based approaches take an alternative route by aligning teacher-student model outputs when evaluating distilled datasets. Notable examples include SRe2L (Yin et al., 2023) and RDED (Sun et al., 2023), where the student model is trained by aligning outputs with outputs of a teacher model on the same batch of synthetic data, specifically by minimizing the Kullback-Leibler (KL) divergence between the student’s predictions and the teacher’s output. In our work, we adopt knowledge distillation as a validation strategy for fair comparisons.
@@
DD research has mainly focused on simpler datasets such as CIFAR (Krizhevsky, 2009), TinyImageNet (Le & Yang, 2015), and DC-BENCH (Cui et al., 2022b). These datasets contain a high proportion of class-specific information, enabling DD methods to extract and synthesize discriminative features more easily. However, research in more complex scenarios has been limited. To address this, we propose the Comp-DD benchmark, which systematically explores dataset distillation complexity by curating subsets from ImageNet-1K with varying degrees of difficulty. This benchmark provides a more rigorous evaluation framework, facilitating deeper exploration of DD in complex, real-world settings and encouraging further advances in the field.
@@ 7 Conclusion
Limitations and Future Work.
EDF dynamically updates Grad-CAM activation maps of synthetic images according to an update frequency. This may introduce extra computation, especially when the IPC is large. Also, we only use Grad-CAM to evaluate discriminative areas of an image in this work. In the future, other indicators that can identify discriminative features of an image can be used jointly to include more perspectives.
Acknowledgement.
We thank Bo Zhao, Guang Li, Dai Liu, Zhenghao Zhao, Songhua Liu, Ruonan Yu, Zhiyuan Liang, and Gongfan Fang for valuable discussions and feedbacks.
Appendix
We organize our appendix as follows.
@@ Acknowledgement.
  • Appendix A: Pseudo code of EDF with detailed explanation.
@@
  • Appendix B.1: Training recipe.
  • Appendix B.2: Evaluation recipe.
  • Appendix B.3: Computing resources required for different settings.
@@
  • Appendix C.1: Cross-architecture evaluation.
  • Appendix C.2: Results of distilled datasets without knowledge-distillation-based evaluation.
  • Appendix C.3: Distorted synthetic images under excessive enhancement factors.
  • Appendix C.4: The changing trend of discriminative areas in EDF distillation.
@@
  • Appendix D.1: Subset details of the Comp-DD benchmark.
  • Appendix D.3: Hyper-parameters of the Comp-DD benchmark.
  • Appendix D.2: More clarifications on the complexity metrics.
@@
  • Appendix E: Visualization of EDF distilled images.
@@
  • Appendix F: More related work of dataset distillation.
Appendix A Algorithm of EDF
@@
1:  Input: $D_{real}$: The real dataset
2:  Input: $D_{syn}$: The synthetic dataset
3:  Input: $\mathcal{A}$: A trajectory-matching based algorithm
4:  Input: $\mathcal{G}$: Grad-CAM model
5:  Input: $K$: Activation maps update frequency
6:  Input: $\alpha$: Threshold of supervision dropout
7:  Input: $T$: Total distillation steps
8:  Input: $\beta$: Enhancement factor
9:  Input: $\mathcal{F}$: Activation map processing function
10:  Input: $r$: Learning rate of synthetic dataset
11:  for $t$ in $10\dots T-$ 1 do
12:     $L\leftarrow\mathcal{A}(D_{syn},D_{real})$                               $\triangleright$ Compute the array of trajectory matching losses
13:     $L^{\prime}\leftarrow Sort(L)$                                            $\triangleright$ Sort $L$ to get ordered losses
14:     $L_{edf}\leftarrow\sum_{i=\alpha|L|}^{|L|}L^{\prime}_{i}$                                 $\triangleright$ Dropout low-loss supervision
15:     if $t\bmod K=0$ then
16:        $M\leftarrow\mathcal{G}(D_{syn})$                                     $\triangleright$ Update activation maps of current $S$
17:     end if
18:     $(\nabla D_{syn})_{EDF}\leftarrow\nabla D_{syn}\circ\mathcal{F}(M,\beta)$      $\triangleright$ Process synthetic image gradients
19:     $D_{syn}\leftarrow D_{syn}-r\cdot(\nabla D_{syn})_{EDF}$          $\triangleright$ Biased update towards discriminative areas
20:  end for
21:  Return $D_{syn}$
@@
Appendix B Experimental Settings
B.1 Training Details
We follow previous trajectory matching works (Du et al., 2022; Guo et al., 2024; Li et al., 2024) to train expert trajectories for one hundred epochs. Hyper-parameters are directly adopted without modification. For distillation, we implement EDF based on DATM (Guo et al., 2024) and PAD (Li et al., 2024), which simultaneously distills soft labels along with images.
We use torch-cam (Fernandez, 2020) for Grad-CAM implementation. Hyper-parameters are listed in Table 9.
B.2 Evaluation Details
To achieve a fair comparison, when comparing EDF with DD methods, we only adopt the set of differentiable augmentations commonly used in previous studies (Zhao & Bilen, 2021b; a; Cazenavette et al., 2022) to train a surrogate model on distilled data and labels.
@@ B.2 Evaluation Details
B.3 Computing Resources
Experiments on IPC 1/10 can be run with 4x Nvidia-A100 80GB GPUs, and experiments on IPC 50 can be run with 8x Nvidia-A100 80GB GPUs. The GPU memory demand is primarily dictated by the volume of synthetic data per batch and the total training iterations the augmentation model undergoes with that data. When IPC becomes large, GPU usage can be optimized by either adopting techniques like TESLA (Cui et al., 2022a) or by scaling down the number of training iterations (”syn_steps”) or shrinking the synthetic data batch size (”batch_syn”).
@@ B.3 Computing Resources
Hyper-parameters $\alpha$ $\beta$ $K$ $T$ batch_syn lr_pixel lr_label syn_steps
ImageNette 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 1 100 250 1000 2.0 40
50 0.375 1 200 100 100 5.0 80
ImageWoof 1 0 1 50 10000 1000 10000 2.0 40
10 0.25 2 100 250 1000 2.0 40
50 0.375 2 200 100 100 5.0 80
ImageMeow 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 250 1000 2.0 40
50 0.375 2 200 200 100 5.0 40
ImageYellow 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 250 1000 3.0 40
50 0.375 1 200 200 100 5.0 40
ImageFruit 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 250 1000 2.0 40
50 0.375 1 200 200 100 5.0 40
ImageSquawk 1 0 1 50 10000 1000 10000 3.0 40
10 0.25 1 100 250 1000 3.0 40
50 0.375 2 200 100 100 5.0 80
Table 9: Hyper-parameters of experiments on ImageNet-1K and nette, woof, meow, fruit, yellow, squawk subsets.
Appendix C Additional experiment results and findings
C.1 Cross-architecture Evaluation
Generalizability on different model architectures is one key property of a well-distilled dataset. To show that EDF can generalize well on different models, we evaluate synthetic images under IPC 10 and 50 of the ImageSquawk subset, on three other standard models, AlexNet (Krizhevsky et al., 2012), VGG11 (Simonyan & Zisserman, 2014), and ResNet18 (He et al., 2015). As shown in Table 12, our distilled datasets outperform random selection and two baseline methods on both IPC10 and IPC50. Compared with IPC10, distilled images under IPC50 can achieve better performance on unseen neural networks. This suggests that EDF’s distillation results have decent generalizability across different architectures, especially when the compressing ratio is smaller which allows distilled datasets to accommodate more discriminative information.

### v2

@@ 4.3 Ablation Study
In this section, we perform ablation studies to analyze key components of our approach. Unless otherwise specified, the following results are all based on ConvNetD5.
@@
We validate the contribution of Discriminative Area Enhancement (DAE) and Common Pattern Dropout (CPD), respectively. As illustrated in Table 5, both DAE and CPD significantly improve the baseline performance. DAE’s biased updates toward high-activation areas effectively enhance the discriminative features in synthetic images. CPD, on the other hand, mitigates the negative influence of common patterns by filtering out low-loss supervision, ensuring that the synthetic images retain their discriminative properties.
DAE CPD ImageWoof ImageMeow ImageYellow
39.2 48.9 65.1
40.3 49.5 66.2
41.1 51.2 67.5
41.8 52.6 68.2
@@
The Discriminative Area Enhancement (DAE) component involves two key factors: the enhancement factor $\beta$ and the threshold for activation maps. Ablation studies (Table 8(a)) show that the best performance is achieved when $\beta$ is between 1 and 2. When $\beta<1$, some discriminative areas are diminished rather than enhanced, as their gradient weights become $<1$. Conversely, excessively large $\beta$ values ($\geq 10$) lead to overemphasis on certain areas, distorting the overall learning process (see Appendix 11.3 for examples of this distortion). Therefore, $\beta$ should be reasonably controlled to balance the emphasis on discriminative regions.
Dynamic thresholding using mean activation values outperforms fixed thresholds (Table 8(b)). This is because the mean adapts to the evolving activation maps during training, whereas a fixed threshold risks either emphasizing low-activation areas if set too low or omitting key discriminative features if set too high.
@@
Table 8: (a) Ablation of the enhancement factor on ImageMeow and ImageWoof. (b) Ablation of the activation threshold on ImageMeow.
@@ 5 Analysis and Discussion
Figure 5: (a) Comparison between high-loss and low-loss supervision distilled images. (b) Comparison of discriminative areas in images produced by initialization, DATM, and EDF. Figures at the bottom are increments made by EDF over the initial image.
@@
Our Discriminative Area Enhancement (DAE) module aims to amplify updates in high-activation areas of synthetic images, as identified by Grad-CAM. To show how DAE enhances discriminative areas, we visualize the same group of images under initialization, DATM distillation, and EDF distillation in Figure 5(b) We also report discriminative area statistics, computed by the percentage of pixels whose activation values are higher than the mean, on each image at the bottom left. As can be discovered, DATM is capable of increasing discriminative regions, while EDF can achieve a more significant enhancement. Visually, the enhancement manifests through an increased number of core objects and enlarged areas of class-specific features. Moreover, EDF’s enhancement is more pronounced, especially when the image has smaller discriminative areas initially, e.g., discriminative features of the first column image increase by 24.5%. These phenomena demonstrate that EDF is effective in capturing and emphasizing discriminative features.
@@
We investigate the impact of the following dropout strategies: (i) dynamic dropout, which includes random selection from all layers, and (ii) static dropout, which includes uniform selection across layers and fixed selection from the first, middle, or last layers. As shown in Table 9, all strategies except EDF’s loss-based dropout lead to performance degradation, with uniform selection and last-layer dropout causing the most significant degradation.
The reasons for this are twofold. First, supervision of low losses—primarily located in the shallow layers of the model—is the main source of common patterns. Discarding supervision from deep layers, where loss values are higher (random selection, uniform selection, or last-layer dropout), reduces the presence of discriminative features. Second, static dropout fails to account for the dynamic nature of low-loss supervision, as trajectory-matching losses vary across layers as the distillation process evolves.
@@ 6 Related Work
DD aims to create compact datasets that maintain performance levels comparable to full-scale datasets. Approaches in DD can be categorized into two primary approaches: matching-based and knowledge-distillation-based.
1) Matching-based approaches are foundational in DD research, focusing on aligning synthetic data with real datasets by capturing essential patterns. Landmark works like gradient matching (DC) [71], distribution matching (DM) [68], and trajectory matching (MTT) [1] extract critical metrics from real datasets, then replicate these metrics in synthetic data. Subsequent research has refined these methods, improving the fidelity of distilled datasets [69, 51, 72, 23, 28, 30, 2, 41, 18]. Data selection techniques have been integrated to synthesize more representative samples [59, 49, 24]. Recent advancements optimize distillation for different image-per-class (IPC) settings, balancing dataset size and information retention [9, 3, 13, 25, 24]. Moreover, soft labels have been widely applied to improve the performance [47, 6, 40, 65]. Current matching-based methods treat pixels uniformly, overlooking discriminative features and supervision signals, limiting their effectiveness on complex datasets.
2) Knowledge-distillation-based approaches take an alternative route by aligning teacher-student model outputs when evaluating distilled datasets. Notable examples include SRe2L [64] and RDED [48]. In our work, we adopt knowledge distillation as a validation strategy for fair comparisons.
@@
Some benchmarks for dataset distillation, such as DC-BENCH [7] and DD-Ranking [26], have been proposed in light of the unfairness issue in previous DD evaluation and provide a more fair comparison scheme. In this work, we propose the Comp-DD benchmark for DD in complex scenarios and offer a new angle to evaluate existing methods. Comp-DD systematically explores dataset distillation complexity by curating subsets from ImageNet-1K with varying degrees of difficulty. This benchmark provides a more rigorous evaluation framework, facilitating deeper exploration of DD in complex, real-world settings and encouraging further advances in the field.
@@ 7 Conclusion
8 Acknowledgement
We thank Bo Zhao, Guang Li, Dai Liu, Zhenghao Zhao, Songhua Liu, Ruonan Yu, Zhiyuan Liang, and Gongfan Fang for valuable discussions and feedback. This work also was supported by the National Research Foundation, Singapore, under its AI Singapore Programme (AISG Award No: AISG2-PhD-2021-08-008). Yang You’s group is being sponsored by NUS startup grant (Presidential Young Professorship), Singapore MOE Tier-1 grant, ByteDance grant, ARCTIC grant, SMI grant (WBS number: A8001104-00-00), Alibaba grant, and Google grant for TPU usage.
Supplementary Material
We organize our supplementary material as follows:
@@ 8 Acknowledgement
  • Appendix 9: Pseudo code of EDF with detailed explanation.
@@
  • Appendix 10.1: Training recipe.
  • Appendix 10.2: Evaluation recipe.
  • Appendix 10.3: Computing resources required for different settings.
@@
  • Appendix 11.1: Cross-architecture evaluation.
  • Appendix 11.2: Results of distilled datasets without knowledge-distillation-based evaluation.
  • Appendix 11.3: Distorted synthetic images under excessive enhancement factors.
  • Appendix 11.4: The changing trend of discriminative areas in EDF distillation.
  • Appendix 11.5: The impact of using different models to extract activation maps.
@@
  • Appendix 12.1: Subset details of the Comp-DD benchmark.
  • Appendix 12.3: Hyper-parameters of the Comp-DD benchmark.
  • Appendix 12.2: More clarifications on the complexity metrics.
@@
  • Appendix 13: Visualization of EDF distilled images.
@@
  • Appendix 14: More related work of dataset distillation.
9 Algorithm of EDF
@@
Input: $D_{real}$: The real dataset
Input: $D_{syn}$: The synthetic dataset
Input: $\mathcal{A}$: A trajectory-matching based algorithm
Input: $\mathcal{G}$: Grad-CAM model
Input: $K$: Activation maps update frequency
Input: $\alpha$: Threshold of supervision dropout
Input: $T$: Total distillation steps
Input: $\beta$: Enhancement factor
Input: $\mathcal{F}$: Activation map processing function
Input: $r$: Learning rate of synthetic dataset
1:   for $t$ in $10\dots T-$ 1 do
2:      $L\leftarrow\mathcal{A}(D_{syn},D_{real})$                               $\triangleright$ Compute the array of trajectory matching losses
3:      $L^{\prime}\leftarrow Sort(L)$                                            $\triangleright$ Sort $L$ to get ordered losses
4:      $L_{edf}\leftarrow\sum_{i=\alpha|L|}^{|L|}L^{\prime}_{i}$                                 $\triangleright$ Dropout low-loss supervision
5:      if $t\bmod K=0$ then
6:         $M\leftarrow\mathcal{G}(D_{syn})$                                     $\triangleright$ Update activation maps of current $S$
7:      end if
8:      $(\nabla D_{syn})_{EDF}\leftarrow\nabla D_{syn}\circ\mathcal{F}(M,\beta)$      $\triangleright$ Process synthetic image gradients
9:      $D_{syn}\leftarrow D_{syn}-r\cdot(\nabla D_{syn})_{EDF}$          $\triangleright$ Biased update towards discriminative areas
10:   end for
11:   Return $D_{syn}$
@@
10 Experimental Settings
10.1 Training Details
We follow previous trajectory matching works [8, 13, 25] to train expert trajectories for one hundred epochs. Hyper-parameters are directly adopted without modification. For distillation, we implement EDF based on DATM [13] and PAD [25], which simultaneously distills soft labels along with images.
We use torch-cam [11] for Grad-CAM implementation. Hyper-parameters are listed in Table 10.
10.2 Evaluation Details
To achieve a fair comparison, when comparing EDF with DD methods, we only adopt the set of differentiable augmentations commonly used in previous studies [69, 68, 1] to train a surrogate model on distilled data and labels.
@@ 10.2 Evaluation Details
10.3 Computing Resources
Experiments on IPC 1/10 can be run with 4x Nvidia-A100 80GB GPUs, and experiments on IPC 50 can be run with 8x Nvidia-A100 80GB GPUs. The GPU memory demand is primarily dictated by the volume of synthetic data per batch and the total training iterations the augmentation model undergoes with that data. When IPC becomes large, GPU usage can be optimized by either adopting techniques like TESLA [6] or by scaling down the number of training iterations (”syn_steps”) or shrinking the synthetic data batch size (”batch_syn”).
@@ 10.3 Computing Resources
Hyper-parameters $\alpha$ $\beta$ $K$ $T$ batch_syn lr_pixel syn_steps
ImageNette 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 1 200 100 100 80
ImageWoof 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 2 200 100 100 80
ImageMeow 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 2 200 200 100 40
ImageYellow 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 1 200 200 100 40
ImageFruit 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 1 200 200 100 40
ImageSquawk 1 0 1 50 10000 1000 10000 40
10 0.25 1 100 250 1000 40
50 0.375 2 200 100 100 80
Table 10: Hyper-parameters of experiments on ImageNet nette, woof, meow, fruit, yellow, squawk subsets.
11 Additional experiment results and findings
11.1 Cross-architecture Evaluation
Generalizability on different model architectures is one key property of a well-distilled dataset. To show that EDF can generalize well on different models, we evaluate synthetic images under IPC 10 and 50 of the ImageSquawk subset, on three other standard models, AlexNet [20], VGG11 [45], and ResNet18 [14]. As shown in Table 15, our distilled datasets outperform random selection and two baseline methods on both IPC10 and IPC50. Compared with IPC10, distilled images under IPC50 can achieve better performance on unseen neural networks. This suggests that EDF’s distillation results have decent generalizability across different architectures, especially when the compressing ratio is smaller which allows distilled datasets to accommodate more discriminative information.